### v1

@@ 5.1 Harsanyi’s Linear Tracing Procedure in Sequence Form
In order to overcome the non-differentiability induced by the boundary constraint conditions when $t\in(0,1]$, we implement a variable substitution. For $x=(x^{i}(w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
$\begin{array}[]{l}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\left\{\begin{array}[]{ll}\psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\in D_{i},\\ (1-\theta(t))x^{i}(w^{i}_{I^{j}_{i}}a)+\theta(t)\psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\notin D_{i},\end{array}\right.\\ \lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\left\{\begin{array}[]{ll}\psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\in D_{i},\\ \theta(t)\psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})&(j,a)\notin D_{i},\end{array}\right.\end{array}$
@@
and $\gamma^{i}(x,t;\emptyset)=1$. This setup follows (17), which avoids the redundant constraints generated by $i\in N,(j,a)\notin D_{i}$ when $t\in(0,1]$. Consequently, we observe that $\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)$ holds for $i\in N,(j,a)\in D_{i}$ and $\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)=x^{i}(w^{i}_{I^{j}_{i}}a),\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=0$ for $i\in N,(j,a)\notin D_{i}$. Substituting $\gamma(x,t)$ and $\lambda(x,t)$ into the system (21) for $\gamma(t)$ and $\lambda$ and subtracting the expression $c(t)(1-\theta(t))\alpha$ from the obtained system, we reach the system (22),
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(x,t),\rho(t)))+\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)\\ \hskip 48.36958pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)-c(t)(1-\theta(t))\alpha(w^{i}_{I^{j}_{i}}a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(x,t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i}.\end{array}$ (22)
This system admits a unique solution at $t=2$, denoted as $(x^{*}(2),\nu^{*}(2))$. The component values are given by $x^{*i}(2;w^{i}_{I^{j}_{i}}a)=\gamma^{0i}(w^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}}-1\text{ for }i\in N,j\in M_{i},a\in A(I^{j}_{i})$, and $\nu^{*i}_{I^{j}_{i}}(2)=1\text{ for }i\in N,j\in M_{i}$.
@@ 5.2 Harsanyi’s Logarithmic Tracing Procedure in Sequence Form
Let $\varepsilon_{0}$ be a positive constant and $\delta=(\delta^{i}(w^{i}):i\in N,w^{i}\in W^{i})$ denote the centroid realization plan profile with $\delta^{i}(w^{i}_{I^{j}_{i}}a)=\delta^{i}(w^{i}_{I^{j}_{i}})/|A(I^{j}_{i})|$. By expanding the influence interval of the logarithmic term to $(0,2]$ in (20), we derive a new artificial game $\Gamma_{s}^{h}(t)$, in which each player $i\in N$ finds their optimal strategy by solving the convex optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-\theta(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)g^{i}(w^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},\rho(t)))\\ &+\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i}(w^{i}_{I^{j}_{i}}a))\ln\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (23)
When $t\in(0,1]$, each player $i$ is incentivized to adjust their strategy closer to the centroid strategy $\delta^{i}$, and (23) increasingly approximates (20) as $\varepsilon_{0}$ goes to 00. Applying the optimality conditions to problem (23) and enforcing $\hat{\gamma}=\gamma(t)$ produces the following polynomial equilibrium system,
$\begin{array}[]{l}(1-\theta(t))g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(t),\rho(t)))+\lambda^{i}(w^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(t;w^{i}_{I^{j}_{i}}a)\lambda^{i}(w^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i},\\ \hskip 162.18062pt0<\gamma^{i}(t;w^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (24)
For $x=(x^{i}(w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
$\begin{array}[]{l}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\psi_{1}(x^{i}(w^{i}_{I^{j}_{i}}a),(\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i})^{1/\kappa_{0}};1,\kappa_{0}),\\ \lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)=\psi_{2}(x^{i}(w^{i}_{I^{j}_{i}}a),(\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i})^{1/\kappa_{0}};1,\kappa_{0}),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\\ \end{array}$ (25)
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(w^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(x,t),\rho(t)))+\lambda^{i}(x,t;w^{i}_{I^{j}_{i}}a)\\ \hskip 48.36958pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)-c(t)(1-\theta(t))\alpha(w^{i}_{I^{j}_{i}}a)=0,\;i\in N,,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,t;w^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(x,t;w^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(w^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i}.\end{array}$ (26)
@@ 6 Numerical Performance
To achieve these objectives, we employed the predictor-corrector method to numerically trace the smooth paths defined by the systems (16), (22), and (26), respectively referred to as LOGB, HLTP, and HLOG ($\varepsilon_{0}=1$). During the tracing procedure, each iteration comprised a predictor to approximate the next solution and a corrector to refine this approximation for improved accuracy. Detailed implementation guidelines can be found in Allgower and Georg [47] and Eaves and Schmedders [45]. The adopted parameter settings including a predictor step size of $0.05t^{0.3}$ and a corrector accuracy of $0.5t^{0.3}$. The successful termination criterion $t<10^{-4}$ was applied, with failure occurring when the number of iterations or computational time surpassed predefined limits. All computations were conducted on a Windows Server 2016 Standard with an Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz (2 processors) and 128GB of RAM.
@@
Figure 1: An Extensive-Form Game from van Damme [6]
@@
Figure 9: An Extensive-Form Game from Mas-Colell et al. [48]
Figure 10: An Extensive-Form Game from Bonanno [49]
@@
To compare the convergence performance of our methods, we employ two structurally distinct types of random extensive-form games, as shown in Figures 1818. Both game types are parameterized by the number of players ($n$), the maximum historical depth ($\mathcal{L}$), and the number of allowable actions per information set ($\mathcal{A}$). In these games, players act cyclically, with the terminal payoffs determined by random integers uniformly distributed between $-10$ and $10$. A detailed explanation of the two game types is provided below.
@@
  • Type 2: As represented in Figures 18, this structural configuration is commonly found in the literature. For odd-indexed players, each information set consists of a single history. In contrast, for even-indexed players, histories are grouped into the same information set only when they share an identical corresponding sequence. The probability that player 00 chooses each of the available actions is equal, and the total number of actions is fixed at $3$, without loss of generality.
Since the number of players does not directly impact the game size, we set $n=3$ for Type 1 games and $n=4$ for Type 2 games, adjusting the other two parameters to control the game size. To realize a comprehensive comparative analysis of the three path-following methods, $20$ random games with distinct payoffs were generated and solved for each parameter configuration $(\mathcal{L},\mathcal{A})$ in both game types. A randomly generated starting point was employed for all three methods in solving each game, and the parameters of the predictor-corrector algorithm remained consistent throughout the entire experiment.

### v3

@@ 5.1 Harsanyi’s linear tracing procedure in sequence form
In order to overcome the non-differentiability induced by the boundary constraint conditions when $t\in(0,1]$, we implement a variable substitution. For $x=(x^{i}(\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
$\begin{array}[]{l}\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)=\psi_{1}(x^{i}(\varpi^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0})\\ \lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)=\psi_{2}(x^{i}(\varpi^{i}_{I^{j}_{i}}a),\theta(t)^{1/\kappa_{0}};\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}},\kappa_{0}),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\end{array}$
@@
and $\gamma^{i}(x,t;\emptyset)=1$. Consequently, we observe that $\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)$ holds for $i\in N,j\in M_{i},a\in A(I^{j}_{i})$. Substituting $\gamma(x,t)$ and $\lambda(x,t)$ into the system (20) for $\gamma(t)$ and $\lambda$ and subtracting the expression $c(t)(1-\theta(t))\alpha$ from the obtained system, we reach the system (21),
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(x,t),\rho(t)))+\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)\\ \hskip 11.38092pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)-c(t)(1-\theta(t))\alpha(\varpi^{i}_{I^{j}_{i}}a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i}.\end{array}$ (21)
At $t=2$, the system possesses a unique solution represented by $(x^{*}(2),\nu^{*}(2))$. The components are expressed as $x^{*i}(2;\varpi^{i}_{I^{j}_{i}}a)=\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)^{1/\kappa_{0}}-1\text{ for }i\in N,j\in M_{i},a\in A(I^{j}_{i})$, and $\nu^{*i}_{I^{j}_{i}}(2)=1\text{ for }i\in N,j\in M_{i}$.
@@ 5.2 Harsanyi’s logarithmic tracing procedure in sequence form
Let $\varepsilon_{0}$ be a positive constant and $\delta=(\delta^{i}(\varpi^{i}):i\in N,\varpi^{i}\in W^{i})$ denote the centroid realization plan profile with $\delta^{i}(\varpi^{i}_{I^{j}_{i}}a)=\delta^{i}(\varpi^{i}_{I^{j}_{i}})/|A(I^{j}_{i})|$. By expanding the influence interval of the logarithmic term to $(0,2]$ in the problem (19), we derive a new artificial game $\Gamma_{s}^{h}(t)$, in which each player $i\in N$ finds their optimal strategy by solving the convex optimization problem,
$\begin{array}[]{rl}\max\limits_{\gamma^{i}(t)}&(1-\theta(t))\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\hat{\gamma}^{-i},\rho(t)))\\ &+\sum\limits_{j\in M_{i}}\sum\limits_{a\in A(I^{j}_{i})}(\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i}(\varpi^{i}_{I^{j}_{i}}a))\ln\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\\ \text{s.t.}&\sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;j\in M_{i}.\end{array}$ (22)
When $t\in(0,1]$, each player $i$ is incentivized to adjust their strategy closer to the centroid strategy $\delta^{i}$, and (22) increasingly approximates (19) as $\varepsilon_{0}$ goes to 0. Applying the optimality conditions to the problem (22) and enforcing $\hat{\gamma}=\gamma(t)$ produces the following polynomial equilibrium system,
$\begin{array}[]{l}(1-\theta(t))g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(t),\rho(t)))+\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)-\nu^{i}_{I^{j}_{i}}\\ \hskip 99.58464pt+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)=0,\;i\in N,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i},\\ \gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a)\lambda^{i}(\varpi^{i}_{I^{j}_{i}}a)=\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}}a)+c(t)(1-\theta(t))\varepsilon_{0}\delta^{i}(\varpi^{i}_{I^{j}_{i}}a),\\ \hskip 85.35826pt0<\gamma^{i}(t;\varpi^{i}_{I^{j}_{i}}a),\;i\in N,j\in M_{i},a\in A(I^{j}_{i}).\end{array}$ (23)
For $x=(x^{i}(\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))\in\mathbb{R}^{n_{0}}$, we define $\gamma(x,t)=(\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$ and $\lambda(x,t)=(\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a):i\in N,j\in M_{i},a\in A(I^{j}_{i}))$, where
[…]
@@
$\begin{array}[]{l}(1-\theta(t))g^{i}(\varpi^{i}_{I^{j}_{i}}a,y(\gamma^{-i}(x,t),\rho(t)))+\lambda^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)\\ \hskip 17.07182pt-\nu^{i}_{I^{j}_{i}}+(1-\theta(t))\zeta^{i}_{I^{j}_{i}}(a)-c(t)(1-\theta(t))\alpha(\varpi^{i}_{I^{j}_{i}}a)=0,\;i\in N,,j\in M_{i},a\in A(I^{j}_{i}),\\ \sum\limits_{a\in A(I^{j}_{i})}\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}}a)-(1-\theta(t))\gamma^{i}(x,t;\varpi^{i}_{I^{j}_{i}})-\theta(t)\gamma^{0i}(\varpi^{i}_{I^{j}_{i}})=0,\;i\in N,j\in M_{i}.\end{array}$ (25)
@@ 6 Numerical performance
To achieve these objectives, we employed the predictor-corrector method to numerically trace the smooth paths defined by the systems (16), (21), and (25), respectively referred to as LOGB, HLTP, and HLOG ($\varepsilon_{0}=1$). In addition, to enhance the smoothness of the paths near $t=1$, we adopted the variable-substitution adjustment technique proposed in [houSequenceformDifferentiablePathfollowing2025]. During the tracing procedure, each iteration comprised a predictor to approximate the next solution and a corrector to refine this approximation for improved accuracy. Detailed implementation guidelines can be found in Allgower and Georg [AllgowerNumericalContinuationMethods1990] and Rheinboldt [rheinboldtNumericalContinuationMethods2000]. The adopted parameter settings including a predictor step size of $0.05t^{0.3}$ and a corrector accuracy of $0.5t^{0.3}$. The successful termination criterion $t<10^{-4}$ was applied, with failure occurring when the number of iterations or computational time surpassed predefined limits. All computations were conducted on a Windows Server 2016 Standard with an Intel(R) Xeon(R) CPU E5-2650 v4 @ 2.20GHz (2 processors) and 128GB of RAM.
@@
Figure 1: An Extensive-Form Game from van Damme [vandammeRelationPerfectEquilibria1984]
@@
Figure 9: An Extensive-Form Game from Mas-Colell et al. [andreuMicroeconomicTheory1995]
Figure 10: An Extensive-Form Game from Bonanno [bonanno2015game]
@@
To compare the convergence performance of our methods, we employ two structurally distinct types of random extensive-form games, as shown in Figures 1818. Both game types are parameterized by the number of players ($n$), the maximum historical depth ($\mathcal{L}$), and the number of allowable actions per information set ($\mathcal{A}$). In these games, players act cyclically, with the terminal payoffs determined by random integers uniformly distributed between 10-10 and $10$. A detailed explanation of the two game types is provided below.
@@
  • Type 2: As represented in Figures 18, this structural configuration is commonly found in the literature. For odd-indexed players, each information set consists of a single history. In contrast, for even-indexed players, histories are grouped into the same information set only when they share an identical corresponding sequence. The probability that player 0 chooses each of the available actions is equal, and the total number of actions is fixed at $3$, without loss of generality.
Since the number of players does not directly impact the game size, we set $n=3$ for Type 1 games and $n=4$ for Type 2 games, adjusting the other two parameters to control the game size. To conduct a thorough comparative evaluation of the three path-following methods, $20$ random games with distinct payoffs were generated and solved for each parameter configuration $(\mathcal{L},\mathcal{A})$ across both types of games. For each game, all three methods utilized the same randomly generated starting point, and the parameters of the predictor-corrector algorithm remained consistent throughout the entire experiment.